A hydrogen atom is made up of a single proton and a single electron. The proton has a charge of +e and the electron has +e . In the “ground state” of the atom, the electron orbits the proton at most probable distance of 5.39 × 10−11 m. Determine the electric force on the electron due to the proton.


1
Expert's answer
2020-05-12T09:59:50-0400

The electric force magnitude is


"F=k\frac{q_1q_2}{r^2}=k\frac{e^2}{r^2}=k\frac{e^2}{r^2}=\\
\space\\=(9\cdot10^9)\frac{(1.602×10^{-19})^2}{(5.39×10^{−11})^2}=7.95\cdot10^{-8}\text{ N}."